# Rethinking Creativity Evaluation: A Critical Analysis of Existing Creativity Evaluations

### v2

@@ Appendix H Dataset examples
Creative writing, Non-creative Fortune’s Folly is a heartwarming comedy that follows the misadventures of Lily Thompson, a young woman who seems to have been born under an unlucky star. Lily, an awkward, parentless young lady, has been evicted from her care home, leaving her junior companion, Daisy, who is eagerly awaiting adoption, heartbroken. Lily’s life has been a series of unfortunate events, from slipping on banana peels to accidentally causing minor explosions in chemistry class. Her bad luck is legendary, and she’s become the town’s favorite source of amusement. However, Lily’s misfortune takes a turn when she stumbles upon a mysterious old woman who claims to be a fortune teller. The fortune teller, Madame Zara, promises to change Lily’s luck if she completes a series of bizarre tasks. With nothing to lose, Lily embarks on a hilarious journey, which includes wrestling a pig, singing opera in a chicken suit, and even attempting to steal the mayor’s prized gnome collection. Meanwhile, Daisy, missing her friend, decides to take matters into her own hands. She embarks on a mission to find a family willing to adopt both her and Lily. Daisy’s attempts to find a family are equally hilarious, involving a series of disastrous interviews with potential parents, including a couple of circus performers, a pair of overly enthusiastic survivalists, and a family of competitive eaters. As Lily completes each task, her luck seems to improve, leading to a series of comedic twists and turns. She accidentally discovers a hidden treasure, saves the town from a disastrous flood, and even manages to win the heart of the handsome local baker, much to the town’s surprise. In the end, Lily’s journey of self-discovery and Daisy’s relentless pursuit of a family lead them to a surprising revelation. The fortune teller, Madame Zara, is actually a wealthy recluse looking for a family to leave her fortune to. Touched by Lily and Daisy’s bond and resilience, she decides to adopt them both, providing them with the family they’ve always wanted. "Fortune’s Folly" is a comedy of errors that proves that sometimes, the greatest fortunes come in the most unexpected ways. It’s a story of friendship, resilience, and the hilarious lengths one will go to change their luck.
@@
Problem-solving, Non-creative You’re outside in the rain and realize you have important papers in your backpack that need to stay dry. You can see water is already beginning to seep through your backpack. Available tools: pencil, umbrella (doesn’t open properly, limited functionality), plastic bag, carabiner, credit card, big paperclip (rusted), old newspaper, shoe lace (frayed, unreliable). How can you protect your important documents from getting wet? To safeguard your documents from moisture: 1. Extract all papers from your backpack. 2. Insert the documents into a waterproof plastic container (bag). 3. Secure the container by creating multiple folds at the top opening to ensure a watertight seal. 4. Reinsert the now-protected paperwork into your backpack for transportation. This method keeps your important documents dry by creating a waterproof barrier between them and any potential moisture that might enter your backpack.
@@
Research ideation, Non-creative Euclidean clustering is a fundamental task with applications in robotics, computer vision, environmental modeling, and many other fields. In this setting, the goal is to identify groups of data points that are close to one another in Euclidean space under a standard distance metric. Traditional methods—such as variants of Lloyd’s algorithm or spectral clustering—approach the problem by first selecting initial cluster centroids and then repeating two main steps: assigning each data point to its nearest centroid and recalculating the centroids as the mean of the points associated with each cluster. This iterative process continues until the centroids stabilize or a fixed number of iterations is reached, with the objective of minimizing the overall sum of squared distances between data points and their centroids. A known limitation of these approaches is that their performance heavily depends on the initialization and may become trapped in local minima. In our work, we derive a closed-form solution to the clustering problem that forgoes the need for such iterative, initialization-sensitive procedures. Instead, we leverage the fact that clustered data tend to lie near a specific subspace in the feature space. The projection operator onto that subspace contains the clustering information in the form of an adjacency matrix, which can be extracted by a simple thresholding operation. Because projection operators are unique, we can determine both the subspace and the corresponding clustering through a straightforward singular value decomposition. Our main theoretical result shows that, when clusters are well defined—meaning they are sufficiently separated and each individual cluster is relatively compact—our closed-form solution is guaranteed to identify the clusters correctly with certainty. In this context, the degree of separation between clusters serves as a proxy for noise, ensuring that the method’s correctness depends only on the level of noise and not on its specific distribution. The proof of this result is both elegant and direct, relying on a few key insights and a classical theorem from matrix perturbation theory. While some practical applications may tolerate heuristic accuracy measures, other sensitive applications demand outputs with certified reliability. For instance, in single-cell sequencing—where identifying correlations between gene activity and cell types is crucial—modern methods often fall short, relying on algorithms that cannot guarantee correctness and on indirect accuracy indicators. Our approach directly addresses this shortcoming. We support our theoretical contributions with extensive experiments on both synthetic data and several real datasets from single-cell sequencing. These experiments not only validate our analysis but also demonstrate that simple adaptations of our closed-form solution yield practical algorithms that gracefully handle increasing noise. In many cases, these variants rival or even exceed the current state-of-the-art in both accuracy and speed. The remainder of the paper is organized as follows. Section 2 outlines the problem and provides an overview of related work. Section 3 presents our main contributions, while Section 4 details the key ideas and proofs behind our theoretical findings. Section 5 describes practical adaptations of our solution, and Section 6 presents our experimental results.

### v3

@@ Appendix H Dataset examples
Creative writing, Uncreative Fortune’s Folly is a heartwarming comedy that follows the misadventures of Lily Thompson, a young woman who seems to have been born under an unlucky star. Lily, an awkward, parentless young lady, has been evicted from her care home, leaving her junior companion, Daisy, who is eagerly awaiting adoption, heartbroken. Lily’s life has been a series of unfortunate events, from slipping on banana peels to accidentally causing minor explosions in chemistry class. Her bad luck is legendary, and she’s become the town’s favorite source of amusement. However, Lily’s misfortune takes a turn when she stumbles upon a mysterious old woman who claims to be a fortune teller. The fortune teller, Madame Zara, promises to change Lily’s luck if she completes a series of bizarre tasks. With nothing to lose, Lily embarks on a hilarious journey, which includes wrestling a pig, singing opera in a chicken suit, and even attempting to steal the mayor’s prized gnome collection. Meanwhile, Daisy, missing her friend, decides to take matters into her own hands. She embarks on a mission to find a family willing to adopt both her and Lily. Daisy’s attempts to find a family are equally hilarious, involving a series of disastrous interviews with potential parents, including a couple of circus performers, a pair of overly enthusiastic survivalists, and a family of competitive eaters. As Lily completes each task, her luck seems to improve, leading to a series of comedic twists and turns. She accidentally discovers a hidden treasure, saves the town from a disastrous flood, and even manages to win the heart of the handsome local baker, much to the town’s surprise. In the end, Lily’s journey of self-discovery and Daisy’s relentless pursuit of a family lead them to a surprising revelation. The fortune teller, Madame Zara, is actually a wealthy recluse looking for a family to leave her fortune to. Touched by Lily and Daisy’s bond and resilience, she decides to adopt them both, providing them with the family they’ve always wanted. "Fortune’s Folly" is a comedy of errors that proves that sometimes, the greatest fortunes come in the most unexpected ways. It’s a story of friendship, resilience, and the hilarious lengths one will go to change their luck.
@@
Problem-solving, Uncreative You’re outside in the rain and realize you have important papers in your backpack that need to stay dry. You can see water is already beginning to seep through your backpack. Available tools: pencil, umbrella (doesn’t open properly, limited functionality), plastic bag, carabiner, credit card, big paperclip (rusted), old newspaper, shoe lace (frayed, unreliable). How can you protect your important documents from getting wet? To safeguard your documents from moisture: 1. Extract all papers from your backpack. 2. Insert the documents into a waterproof plastic container (bag). 3. Secure the container by creating multiple folds at the top opening to ensure a watertight seal. 4. Reinsert the now-protected paperwork into your backpack for transportation. This method keeps your important documents dry by creating a waterproof barrier between them and any potential moisture that might enter your backpack.
@@
Research ideation, Uncreative Euclidean clustering is a fundamental task with applications in robotics, computer vision, environmental modeling, and many other fields. In this setting, the goal is to identify groups of data points that are close to one another in Euclidean space under a standard distance metric. Traditional methods—such as variants of Lloyd’s algorithm or spectral clustering—approach the problem by first selecting initial cluster centroids and then repeating two main steps: assigning each data point to its nearest centroid and recalculating the centroids as the mean of the points associated with each cluster. This iterative process continues until the centroids stabilize or a fixed number of iterations is reached, with the objective of minimizing the overall sum of squared distances between data points and their centroids. A known limitation of these approaches is that their performance heavily depends on the initialization and may become trapped in local minima. In our work, we derive a closed-form solution to the clustering problem that forgoes the need for such iterative, initialization-sensitive procedures. Instead, we leverage the fact that clustered data tend to lie near a specific subspace in the feature space. The projection operator onto that subspace contains the clustering information in the form of an adjacency matrix, which can be extracted by a simple thresholding operation. Because projection operators are unique, we can determine both the subspace and the corresponding clustering through a straightforward singular value decomposition. Our main theoretical result shows that, when clusters are well defined—meaning they are sufficiently separated and each individual cluster is relatively compact—our closed-form solution is guaranteed to identify the clusters correctly with certainty. In this context, the degree of separation between clusters serves as a proxy for noise, ensuring that the method’s correctness depends only on the level of noise and not on its specific distribution. The proof of this result is both elegant and direct, relying on a few key insights and a classical theorem from matrix perturbation theory. While some practical applications may tolerate heuristic accuracy measures, other sensitive applications demand outputs with certified reliability. For instance, in single-cell sequencing—where identifying correlations between gene activity and cell types is crucial—modern methods often fall short, relying on algorithms that cannot guarantee correctness and on indirect accuracy indicators. Our approach directly addresses this shortcoming. We support our theoretical contributions with extensive experiments on both synthetic data and several real datasets from single-cell sequencing. These experiments not only validate our analysis but also demonstrate that simple adaptations of our closed-form solution yield practical algorithms that gracefully handle increasing noise. In many cases, these variants rival or even exceed the current state-of-the-art in both accuracy and speed. The remainder of the paper is organized as follows. Section 2 outlines the problem and provides an overview of related work. Section 3 presents our main contributions, while Section 4 details the key ideas and proofs behind our theoretical findings. Section 5 describes practical adaptations of our solution, and Section 6 presents our experimental results.